Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0
solution:8, only solution

Answer by stanbon(75887)

Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0
solution:8, only solution
--------------
f(x) = (x-8)(x-8)
f(x) = x^2-16x+64
========================
cheers,
Stan H.